o save money, a local charity organization wants to target its mailing requests for donations to individuals who are most supportive of its cause. They ask a sample of 5 men and 5 women to rate the importance of their cause on a scale from 1 (not important at all) to 7 (very important). The ratings for men were M1 = 6.3. The ratings for women were M2 = 5.5. If the estimated standard error for the difference is equal to 0.25, then consider the following. (a) Find the confidence limits at an 80% CI for these two independent samples.
Answer #1

M1 = 6.3, M2 = 5.5, n1 = 5, n2 = 5

se = 0.25

80% Confidence interval :

At α = 0.2 and df = n1+n2-2 = 8, two tailed critical value, t-crit = T.INV.2T(0.2, 8) = 1.397

Lower Bound = (M1 - M2) - t-crit*se = (6.3 - 5.5) - 1.397*0.25 = 0.4508

Upper Bound = (M1 - M2) + t-crit*se = (6.3 - 5.5) + 1.397*0.25 = 1.1492